12. An astronaut experiences a force of gravity on Earth of 800 N. Neptune has 17 times the mass and 3.9 times the radius of Earth. What would the astronaut weigh, in newtons, on Neptune?


1
Expert's answer
2021-10-31T18:13:28-0400

Given:

"F_E=800\:\rm N"

"M_N=17M_E"

"R_N=3.9R_E"


The force of gravity on an astronaut with mass "m" at the surface of planet with mass "M" and radius "R" is given by

"F=G\frac{Mm}{R^2}"

Hence

"\frac{F_N}{F_E}=\frac{M_N}{M_E}*\frac{R_E^2}{R_N^2}"

"\frac{F_N}{F_E}=17*1/3.9^2=1.1"

Finally

"F_N=F_E*1.1=800*1.1=880\:\rm N"